# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v1

@@ Abstract
The burgeoning utilization of Large Language Models (LLMs) in scientific research necessitates advanced benchmarks capable of evaluating their understanding and application of scientific knowledge comprehensively. To address this need, we introduce the SciKnowEval benchmark, a novel framework that systematically evaluates LLMs across five progressive levels of scientific knowledge: studying extensively, inquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. These levels aim to assess the breadth and depth of scientific knowledge in LLMs, including knowledge coverage, inquiry and exploration capabilities, reflection and reasoning abilities, ethic and safety considerations, as well as practice proficiency. Specifically, we take biology and chemistry as the two instances of SciKnowEval and construct a dataset encompassing 50K multi-level scientific problems and solutions. By leveraging this dataset, we benchmark 20 leading open-source and proprietary LLMs using zero-shot and few-shot prompting strategies. The results reveal that despite achieving state-of-the-art performance, the proprietary LLMs still have considerable room for improvement, particularly in addressing scientific computations and applications. We anticipate that SciKnowEval will establish a comprehensive standard for benchmarking LLMs in science research and discovery, and promote the development of LLMs that integrate scientific knowledge with strong safety awareness. The dataset and code are publicly available at https://github.com/hicai-zju/sciknoweval.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated an impressive capability in storing and recalling world knowledge, continuously expanding the boundaries of artificial intelligence. Their exceptional performance has permeated diverse specialized domains, including the scientific domain, leading to the emergence of scientific LLMs, such as Galactica [33], SciGLM [42], and ChemLLM [43]. To steadily advance scientific research, it is crucial to establish reliable benchmarks that comprehensively demonstrate these models’ understanding and application of scientific knowledge.
@@
In response to these deficiencies, in this study, we adopt a distinctive perspective, “LLMs as Scientists”, to revisit the evaluation in the scientific domain. We draw inspiration from the profound principles of Confucius outlined in the ancient Chinese philosophy “Doctrine of the Mean” , and present a novel Scientific Knowledge Evaluation benchmark, referred to as SciKnowEval, as illustrated in Fig. 1. This benchmark aims to assess LLMs based on their proficiency in five progressive levels: studying extensively, enquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. Each level offers a unique perspective on evaluating the capabilities of LLMs in handling scientific knowledge.
@@
Figure 1: Illustration of SciKnowEval. We consider LLMs as Scientists and evaluate their performance based on a hierarchical framework inspired by the ancient Chinese philosophy from “ Doctrine of the Mean”, encompassing five progressive levels that span from knowledge acquisition to application. We also construct a large-scale dataset comprising various scientific domain tasks at different levels, along with corresponding evaluation criteria, leading to a reliable benchmark of LLMs.
Table 1: Comparison of SciKnowEval with other benchmarks. “SE” refers to Studying Extensively, “EE” refers to Enquiring Earnestly, “TP” refers to Thinking Profoundly, “DC” refers to Discerning Clearly, “PA” refers to Practicing Assiduously. “_sci” represents the subset of scientific questions.
@@
AGIEval_sci Phy., Geo., Bio., Chem., Math ✓ ✓ MCQ Exam 8,062
@@
SciQ Phy., Bio., Chem. ✓ ✓ MCQ ✓ Crowdsourcing 1,000
@@
SciBench Phy., Chem., Math ✓ GEN ✓ Textbook 789
SciEval Phy., Bio., Chem. ✓ ✓ ✓ MCQ, T/F GEN Website, Knowledge base 15,901
@@
SciKnowEval Bio., Chem. ✓ ✓ ✓ ✓ ✓ MCQ, T/F RE, GEN Literature, Self-gen, Textbook, Database 50,048
@@
  • We propose a multi-level scientific knowledge evaluation framework according to the profound principles of Confucius. Each level targets a critical aspect of knowledge handling by LLMs, ranging from basic knowledge understanding to the practical application of knowledge in real-world scenarios.
  • We construct a large-scale evaluation dataset comprised of diverse biological and chemical problems, accompanied by corresponding solutions and rigorous evaluation metrics, facilitating an extensive assessment of the breadth and depth of scientific knowledge encapsulated in LLMs.
  • We evaluate a wide range of advanced LLMs (including 12 general-purpose LLMs and 8 scientific LLMs) and rank their performance with the SciKnowEval dataset, elucidating both their strength and weaknesses.
@@ 2 Related Works
Assessing proficiency in scientific knowledge is a crucial aspect of LLM evaluation. For example, renowned benchmarks like MMLU [15], AGIEval [45], and ARC [7] have incorporated a number of scientific questions into their assessments. Recently, with the rapid application of LLMs in science research, numerous benchmarks tailored for scientific domains have been developed. SciQ [37] encompasses ~13K crowdsourced science examination questions, covering diverse subjects including Physics, Chemistry, Biology, and more. ScienceQA [24] consists of ~21K multimodal scientific questions and answers (QA) collected from elementary and high school science curricula, focusing on evaluating the interpretability of LLMs in addressing scientific problems. SciBench [36] collects open-ended questions from college-level textbooks in physics, chemistry, and mathematics to assess reasoning abilities for complex scientific problems. SciEval [31], based on Bloom’s Taxonomy, offers a multidisciplinary, multi-level dataset to assess understanding, application, computation, and research abilities. SciAssess [4] introduces a multimodal scientific literature analysis dataset, encompassing tasks like information extraction and chart/table QA. SciMT-Safety [14] presents a red-teaming benchmark with 432 malicious queries, focusing on potential misuse risks of LLMs. ChemLLMBench [13] establishes a comprehensive benchmark to evaluate the abilities of LLMs on eight chemistry tasks. Bioinfo-Bench [6] is a simple LLM benchmark for bioinformatics skills evaluation, comprising 200 questions derived from bioinformatics literature. Table 1 has reported the statistics of these benchmarks. One can observe that they suffer from several limitations, such as inadequate scope of ability examination, limited data size, and insufficient task diversity. In this study, we develop a systematic scientific knowledge evaluation framework and construct a large-scale dataset with multi-level domain tasks to benchmark the general and scientific LLMs.
@@ 3.1 Design Philosophy
The design philosophy of SciKnowEval is inspired by the profound principles of Confucius elucidated in the ancient Chinese book “Doctrine of the Mean”: studying extensively, enquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. This principle reflects the five progressive levels in the process from acquiring to applying knowledge. In this study, we regard LLMs as Scientists and utilize this concept to evaluate them. Specifically, each level provides a perspective to assess the proficiency of LLMs, as described below.
  • L1: Studying extensively (i.e., knowledge coverage). This dimension evaluates the breadth of an LLM’s knowledge across various scientific domains. It measures the model’s ability to remember and understand a wide range of scientific concepts.
  • L2: Enquiring earnestly (i.e., knowledge enquiry and exploration). This aspect focuses on the LLM’s capacity for deep enquiry and exploration within scientific contexts, such as analyzing scientific texts, identifying key concepts, and questioning relevant information.
  • L3: Thinking profoundly (i.e., knowledge reflection and reasoning). This criterion examines the model’s capacity for critical thinking, logical deduction, numerical calculation, function prediction, and the ability to engage in reflective reasoning to solve problems.
  • L4: Discerning clearly (i.e., knowledge discernment and safety assessment). This aspect evaluates the LLM’s ability to make correct, secure, and ethical decisions based on scientific knowledge, including assessing the harmfulness and toxicity of information, and understanding the ethical implications and safety concerns related to scientific endeavors.
  • L5: Practicing assiduously (i.e., knowledge practice and application). The final dimension assesses the LLM’s capability to apply scientific knowledge effectively in real-world scenarios, such as analyzing complex scientific problems and creating innovative solutions.
Building upon the above design philosophy, we develop the SciKnowEval benchmark specifically tailored for assessing multi-level scientific knowledge in LLMs. In particular, we undertake meticulous designs in terms of data scale, diversity and quality when constructing the evaluation dataset: 1) Large-scale. We architect our dataset to be large-scale, enabling a more accurate and robust assessment of LLMs. 2) Multi-level. We design and construct our datasets to encompass a wide range of tasks, spanning multiple levels of scientific knowledge, to comprehensively assess the breadth and depth of knowledge in LLMs. 3) High-quality. We prioritize the quality of our data through rigorous quality control measures, ensuring the reliability of the proposed dataset.
@@ 3.2 Data Collection Methods
Literature and textbooks cover a broad range of scientific knowledge, and leveraging this data will facilitate a comprehensive evaluation of LLMs’ capabilities in the biological and chemical sciences. We collect massive papers from article preprint platforms (e.g., BioRxiv), literature databases (e.g., PubMed), and textbook databases (e.g., LibreTexts). We utilize LLMs to automate the procedures of QA pair generation. Specifically, following domain experts’ advice, we carefully design effective prompts for literature QA tasks. These prompts exhibited in Appendix A4 guide the LLM to extract relevant professional knowledge from literature and textbook paragraphs, enabling it to generate new QA pairs around this expertise. To ensure quality assessment of the generated questions, we emphasize in the prompts that answers must be explicitly found in the original text without introducing any external information.
@@
We sample additional biological and chemical QAs from existing open-source scientific benchmarks, including MedMCQA [28], SciEval [31], MMLU [15], XieZhi [12], PubMedQA [17], and HarmfulQA [3]. To mitigate the risk of data contamination and leakage in these benchmarks, we employ LLMs to refactor these QAs in various forms, such as question rewriting and option reordering. Moreover, in cases where some QAs lack explicit annotations indicating their corresponding levels in SciKnowEval, LLMs are utilized to automatically categorize the data into distinct levels.
@@
To enhance the variety and scope of tasks in our dataset, we select several biological and chemical databases and transform them into textual formats suitable for evaluating LLMs. These databases mainly include molecular (e.g., PubChem [19]), protein (e.g., UniProtKB [8]), and cellular-related (e.g., SHARE-seq [25]) sequence information, which contain annotations related to structure, properties, and functions. We can utilize these annotations to construct QA pairs. Specifically, we first conduct preliminary quality screening, such as filtering out chemically invalid SMILES from PubChem using the RDKit library. We then design multiple question templates to transform the structured sequence-annotation pairs into natural language formats, including multiple-choice questions and true/false questions.
@@
3.3 Datasets for Biology and Chemistry
Based on the above data collection methods and the design philosophy of SciKnowEval, we construct a dataset for Biology and Chemistry. Table 2 shows the overview of our dataset, which encompasses a broad spectrum of domain tasks, ensuring a comprehensive evaluation of LLMs. More details of each task and corresponding instances can be found in Appendix A5. The data quality control, validation and copyright compliance are reported in Appendix A6.
Table 2: Overview of the SciKnowEval dataset for Biology and Chemistry. Abbr., MCQ: multiple choice questions; T/F: true/false; CLS: classification; RE: relation extraction; GEN: generative task. Data collection methods I, II and III are in Fig. 2. The detailed data source listed in Appendix A3.
@@ 3.3 Datasets for Biology and Chemistry
Biology L1 Biological Literature QA MCQ Literature Corpus I 14,869
@@
Biomedical Judgment and Interpretation T/F PubMedQA II 904
Compound-Disease Relation Extraction RE Bohrium II 867
@@
Hypothesis Verification T/F LibreTexts I 619
Reasoning and Interpretation MCQ LibreTexts I 647
L3 Solubility Prediction MCQ PEER, DeepSol III 201
$\beta$-lactamase Activity Prediction MCQ PEER, Envision III 209
Fluorescence Prediction MCQ PEER, Sarkisyan’s III 205
GB1 Fitness Prediction MCQ PEER, FLIP III 201
Stability Prediction MCQ PEER, Rocklin’s III 203
Protein-Protein Interaction MCQ STRING, SHS27K, SHS148K III 205
@@
L4 Biological Harmful QA GEN Self-generated I 297
@@
Biological Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 194
L5 Biological Protocol Procedure Design GEN Protocol Journal I 591
Biological Protocol Reagent Design GEN Protocol Journal I 565
Protein Captioning GEN UniProtKB III 937
Protein Design GEN UniProtKB III 860
@@
Chemistry L1 Molecular Name Conversion MCQ PubChem III 1,008
Molecular Property Identification MCQ, T/F MoleculeNet III 1,625
Chemical Literature QA MCQ Literature Corpus I 6,316
L2 Reaction Mechanism Inference MCQ LibreTexts I 269
Compound Identification and Properties MCQ LibreTexts I 497
Doping Extraction RE NERRE II 821
Detailed Understanding MCQ LibreTexts I 626
Text Summary GEN LibreTexts I 692
Hypothesis Verification T/F LibreTexts I 544
Reasoning and Interpretation MCQ LibreTexts I 516
L3 Molar Weight Calculation MCQ PubChem III 1,042
Molecular Property Calculation MCQ MoleculeNet II 740
Molecular Structure Prediction MCQ PubChem III 608
Reaction Prediction MCQ USPTO-Mixed II 1,122
Retrosynthesis MCQ USPTO-50k II 1,122
Balancing Chemical Equation GEN WebQC III 535
Chemical Calculation MCQ XieZhi, SciEval, MMLU II 269
L4 Chemical Harmful QA GEN Proposition-65, ILO III 454
Molecular Toxicity Prediction MCQ, T/F Toxric III 870
Chemical Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 531
L5 Molecular Captioning GEN ChEBI-20 II 943
Molecular Generation GEN ChEBI-20 II 897
Chemical Protocol Procedure Design GEN Protocol Journal I 74
Chemical Protocol Reagent Design GEN Protocol Journal I 129

### v2

@@ Abstract
Large language models (LLMs) have gained increasing prominence in scientific research, but there is a lack of comprehensive benchmarks to fully evaluate their proficiency in understanding and mastering scientific knowledge. To address this need, we introduce the SciKnowEval benchmark, a novel framework that systematically evaluates LLMs across five progressive levels of scientific knowledge: studying extensively, inquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. These levels aim to assess the breadth and depth of scientific knowledge in LLMs, including memory, comprehension, reasoning, discernment, and application. Specifically, we first construct a large-scale evaluation dataset encompassing 70K multi-level scientific problems and solutions in the domains of biology, chemistry, physics, and materials science. By leveraging this dataset, we benchmark 26 advanced open-source and proprietary LLMs using zero-shot and few-shot prompting strategies. The results reveal that despite the state-of-the-art performance of proprietary LLMs, there is still significant room for improvement, particularly in addressing scientific reasoning and applications. We anticipate that SciKnowEval will establish a standard for benchmarking LLMs in science research and promote the development of stronger scientific LLMs. The dataset and code are publicly available at https://scimind.ai/sciknoweval.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated an impressive capability in storing and recalling world knowledge, continuously expanding the boundaries of artificial intelligence. Their exceptional performance has permeated diverse specialized domains, including the scientific domain, leading to the emergence of scientific LLMs, such as Galactica [36], SciGLM [46], and ChemLLM [47]. To steadily advance scientific research, it is crucial to establish reliable benchmarks that comprehensively evaluate these models’ capability in handling scientific knowledge.
@@
In response to these deficiencies, in this study, we adopt a distinctive perspective, “LLMs as Scientists”, to revisit the evaluation in the scientific domain. We draw inspiration from the profound principles of Confucius outlined in the ancient Chinese philosophy “Doctrine of the Mean” , and present a novel Scientific Knowledge Evaluation benchmark, referred to as SciKnowEval, as illustrated in Fig. 1. This benchmark aims to assess LLMs based on their proficiency in five progressive levels: studying extensively, enquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. Each level offers a unique perspective on evaluating the capabilities of LLMs in handling scientific knowledge, including memory, comprehension, reasoning, discernment, and application.
@@
Figure 1: Illustration of SciKnowEval. We consider LLMs as scientists and evaluate their performance based on a hierarchical framework inspired by the ancient Chinese philosophy from “ Doctrine of the Mean”, encompassing five progressive levels that span from knowledge memory to application. We also construct a large-scale dataset comprising various scientific domain tasks at different levels, along with corresponding evaluation criteria, leading to a reliable benchmark of LLMs.
Table 1: Comparison between SciKnowEval and other benchmarks.
@@
AGIEval_sci Phys., Geo., Bio., Chem., Math ✓ ✓ MCQ Exam 8,062
@@
SciQ Phys., Bio., Chem. ✓ ✓ MCQ ✓ Crowdsourcing 1,000
@@
SciBench Phys., Chem., Math ✓ GEN ✓ Textbook 789
SciEval Phys., Bio., Chem. ✓ ✓ ✓ MCQ, T/F GEN Website, Knowledge base 15,901
@@
SciEx Computer Science ✓ ✓ GEN ✓ Exam 154
ChemBench Chem. ✓ ✓ ✓ ✓ MCQ, GEN ✓ Exam, Book Knowledge base 7059
LAB-Bench Bio. ✓ ✓ ✓ MCQ ✓ Literature, Knowledge base 2400
SciKnowEval Bio., Chem., Phys., Mat. ✓ ✓ ✓ ✓ ✓ MCQ, T/F RE, GEN Literature, Textbook, Database 70,203
@@
  • We propose a multi-level scientific knowledge evaluation framework that targets critical aspects of knowledge handling by LLMs, encompassing memory, comprehension, reasoning, discernment, and application.
  • We construct a large-scale evaluation dataset comprised of diverse scientific problems from the domains of biology, chemistry, physics and material science, accompanied by corresponding solutions and evaluation metrics, facilitating an extensive assessment of the breadth and depth of scientific knowledge encapsulated in LLMs.
  • We evaluate a wide range of advanced LLMs (including 18 general-purpose LLMs and 8 scientific LLMs) and rank their performance with the SciKnowEval dataset, elucidating both their strength and weaknesses.
@@ 2 Related Works
Assessing proficiency in scientific knowledge is a crucial aspect of LLM evaluation. For example, renowned benchmarks like MMLU [18], AGIEval [49], and ARC [6] have incorporated a number of scientific questions into their assessments. Recently, with the rapid application of LLMs in science research, numerous benchmarks tailored for scientific domains have been developed. SciQ [40] encompasses 13K crowdsourced science examination questions, covering diverse subjects including Physics, Chemistry, Biology, and more. ScienceQA [27] consists of 21K multimodal scientific questions and answers (QA) collected from elementary and high school science curricula, focusing on evaluating the interpretability of LLMs in addressing scientific problems. SciBench [39] collects open-ended questions from college-level textbooks in physics, chemistry, and mathematics to assess reasoning abilities for complex scientific problems. SciEval [35], based on Bloom’s Taxonomy, offers a multidisciplinary, multi-level dataset to assess understanding, application, computation, and research abilities. SciAssess [4] introduces a multimodal scientific literature analysis dataset, encompassing tasks like information extraction and chart/table QA. SciMT-Safety [17] presents a red-teaming benchmark with 432 malicious queries, focusing on potential misuse risks of LLMs. ChemLLMBench [16] establishes a comprehensive benchmark to evaluate the abilities of LLMs on eight chemistry tasks. Bioinfo-Bench [5] is a simple LLM benchmark for bioinformatics skills evaluation, comprising 200 questions derived from bioinformatics literature. SciEx [8] compiles 154 multilingual and multimodal questions from the exam banks of seven real university computer science courses to evaluate reasoning and application skills. ChemBench [29] establishes a comprehensive benchmark consisting of 7,059 questions designed to assess the expertise and safety of LLMs across 11 subfields of chemistry. LAB-Bench [24] introduces the language agent biology benchmark, which includes over 2,400 multiple-choice questions aimed at evaluating AI systems on a range of practical biology research capabilities.
Table 1 has reported the statistics of these benchmarks. One can observe that they suffer from several limitations, such as inadequate scope of ability examination, limited data size, and insufficient task diversity. In this study, we develop a systematic scientific knowledge evaluation framework and construct a large-scale dataset with multi-level domain tasks to benchmark the general and scientific LLMs.
@@ 3.1 Design Philosophy
The design philosophy of SciKnowEval is inspired by the profound principles of Confucius elucidated in the ancient Chinese book “Doctrine of the Mean”: Studying extensively, Enquiring earnestly, Thinking profoundly, Discerning clearly, and Practicing assiduously. This principle reflects the five progressive levels in the human learning process. In this study, we regard LLMs as Scientists and utilize this concept to evaluate them. Specifically, each level provides a perspective to assess the proficiency of LLMs, as described below.
  • L1: Studying extensively (i.e., knowledge memory). This dimension evaluates the breadth of an LLM’s knowledge across various scientific domains. It measures the model’s ability to remember a wide range of scientific concepts.
  • L2: Enquiring earnestly (i.e., knowledge comprehension). This aspect focuses on the LLM’s capacity for inquiry and exploration within scientific contexts, such as analyzing scientific texts, identifying key concepts, and questioning relevant information.
  • L3: Thinking profoundly (i.e., knowledge reasoning). This criterion examines the model’s capacity for critical thinking, logical deduction, numerical calculation, function prediction, and the ability to engage in reflective reasoning to solve problems.
  • L4: Discerning clearly (i.e., knowledge discernment). This aspect evaluates the LLM’s ability to make correct, secure, and ethical decisions based on scientific knowledge, including assessing the harmfulness and toxicity of information, and understanding the ethical implications and safety concerns related to scientific endeavors.
  • L5: Practicing assiduously (i.e., knowledge application). The final dimension assesses the LLM’s capability to apply scientific knowledge effectively in real-world scenarios, such as solving complex scientific problems and creating innovative solutions.
Building upon the above design philosophy, we develop the SciKnowEval benchmark specifically tailored for assessing multi-level scientific knowledge in LLMs. In particular, we undertake meticulous designs in terms of data scale, diversity and quality when constructing the evaluation dataset:
  • Large-scale. We architect our dataset to be large-scale, enabling a more accurate and robust assessment of LLMs.
  • Multi-level. We design and construct our datasets to encompass a wide range of tasks, spanning multiple levels of scientific knowledge, to comprehensively assess the breadth and depth of knowledge in LLMs.
  • High-quality. We prioritize the quality of our data through rigorous quality control measures, ensuring the reliability of the proposed dataset.
@@ 3.2 Data Collection Methods
Literature and textbooks cover a broad range of scientific knowledge, and leveraging this data will facilitate a comprehensive evaluation of LLMs’ capabilities in the scientific domains. We collect massive papers from article preprint platforms (e.g., BioRxiv), literature databases (e.g., PubMed), and textbook databases (e.g., LibreTexts). We utilize LLMs to automate the procedures of QA pair generation. Specifically, following domain experts’ advice, we carefully design effective prompts for literature QA tasks. These prompts exhibited in Appendix A5 guide the LLM to extract relevant professional knowledge from literature and textbook paragraphs, enabling it to generate new QA pairs around this expertise. To ensure quality assessment of the generated questions, we emphasize in the prompts that answers must be explicitly found in the original text without introducing any external information.
@@
We sample additional QAs from existing open-source scientific benchmarks, including MedMCQA [31], SciEval [35], MMLU [18], XieZhi [15], PubMedQA [19], and HarmfulQA [3]. To mitigate the risk of data contamination and leakage in these benchmarks, we employ LLMs to refactor these QAs in various forms, such as question rewriting and option reordering. Moreover, in cases where some QAs lack explicit annotations indicating their corresponding levels in SciKnowEval, LLMs are utilized to automatically categorize the data into distinct levels.
@@
To enhance the variety and scope of tasks in our dataset, we select several structured databases and transform them into textual formats suitable for evaluating LLMs. These databases mainly include molecular (e.g., PubChem [21]), protein (e.g., UniProtKB [7]), and cellular-related (e.g., SHARE-seq [28]) sequence information, which contain annotations related to structure, properties, and functions. We can utilize these annotations to construct QA pairs. Specifically, we first conduct preliminary quality screening, such as filtering out chemically invalid SMILES from PubChem using the RDKit library. We then design multiple question templates to transform the structured sequence-annotation pairs into natural language formats, including multiple-choice questions, true/false questions, and short-answer questions.
@@
3.3 Data Quality Control
The primary concern of data quality is the QAs generated from literature corpus. To ensure the generated data with high quality, we employed a three-stage data screening process:
  • Initial screening by LLMs. We first explicitly instructed LLM during data generation that the correct options must be clearly identifiable from the provided literature snippets. After data generation, we prompt LLMs to simulate an open-book exam task to determines whether each question’s answer can be found in the corresponding snippet.
  • Human evaluation. We randomly selected approximately 5% of the generated questions and provided them with two domain experts. During the evaluation, we used the instructions in Table A10 to guide the evaluators, asking them to thoroughly assess the data and classify it into binary categories of “Yes” and “No” for quality.
  • Post-screening by LLMs. We employed LLMs to summarize the failure types of the identified low-quality entries, and added them into the prompt to conduct a full dataset quality assessment, discarding similar types of low-quality questions.
More details of data quality control are provided in Appendix A7.
3.4 Final Datasets
Based on the above data collection and screening process, we construct the SciKnowEval dataset, consisting of four subsets for Biology, Chemistry, Physics and Materials, respectively. Table 3.4 shows the overall statistics of the constructed dataset. Table 3 provides an overview of the biological dataset, and Table A8 summarizes the datasets for other domains. Detailed question examples are available in Appendix A6.
Table 2: Statistics of the SciKnowEval dataset
Statistics Number
Total Questions 70,203
Subjects/Tasks 4/78
L1 Questions 39,264 (55.93%)
L2 Questions 12,896 (18.37%)
L3 Questions 8,368 (11.92%)
L4 Questions 5,257 (7.49%)
L5 Questions 4,418 (6.29%)
Statistics Number
Average question length 50.38
Average option length 6.25
Average answer length 56.60
Multiple-choice Questions 52,770 (75.17%)
Constrained Generation Question 10,722 (15.27%)
True or False Questions 4,723 (6.73%)
Relation Extraction Question 1,988 (2.83%)
Table 3: Overview of the biological dataset in SciKnowEval. Abbr., MCQ: multiple choice questions; T/F: true/false; CLS: classification; RE: relation extraction; GEN: generative task. Data collection methods I, II and III are in Fig. 2. The detailed data source listed in Appendix A4. The overview of other domain datasets (chemistry, physics and materials) is presented in Table A8.
@@ 3.4 Final Datasets
Bio. L1 Biological Literature QA MCQ Literature Corpus I 14,862
@@
Biomedical Judgment and Interpretation T/F PubMedQA II 947
Compound-Disease Relation Extraction RE Bohrium II 500
@@
Hypothesis Verification T/F LibreTexts I 618
Reasoning and Interpretation MCQ LibreTexts I 648
L3 Solubility Prediction MCQ PEER, DeepSol III 207
$\beta$-lactamase Activity Prediction MCQ PEER, Envision III 203
Fluorescence Prediction MCQ PEER, Sarkisyan’s III 203
GB1 Fitness Prediction MCQ PEER, FLIP III 208
Stability Prediction MCQ PEER, Rocklin’s III 204
Protein-Protein Interaction MCQ STRING, SHS27K, SHS148K III 207
@@
L4 Biological Harmful QA GEN Website I 297
@@
Biological Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 192
L5 Biological Protocol Procedure Design GEN Protocol Journal I 577
Biological Protocol Reagent Design GEN Protocol Journal I 588
Protein Design GEN UniProtKB III 949